# Can Prompt Difficulty be Online Predicted for Accelerating RL Finetuning of Reasoning Models?

### v4

@@
Abstract
Recent advances have witnessed the effectiveness of reinforcement learning (RL) finetuning in enhancing the reasoning capabilities of large language models (LLMs). The optimization process often requires numerous iterations to achieve satisfactory performance, resulting in high computational costs due to the need for frequent prompt evaluations under intensive LLM interactions and repeated policy updates. Appropriate online prompt selection methods reduce iteration steps by prioritizing informative prompts during training, while the pipeline’s reliance on exhaustive prompt evaluation and subset selection for optimization still incurs substantial computational overhead due to frequent LLM inference calls. Distinguished from these direct evaluate-then-select schemes, this work investigates iterative approximate evaluation for arbitrary prompts and introduces Model Predictive Prompt Selection (MoPPS), a Bayesian risk-predictive framework that online estimates prompt difficulty without requiring costly LLM interactions. Technically, MoPPS models each prompt’s success rate as a latent variable, performs streaming Bayesian inference, and employs posterior sampling in a constructed multi-armed bandit machine, enabling sample efficient and adaptive prompt selection. Extensive experiments across mathematics, planning, and vision-based geometry tasks show that MoPPS reliably predicts prompt difficulty and accelerates training with significantly reduced LLM rollouts. Our code is available at https://github.com/thu-rllab/MoPPS.
Figure 1: Spearman rank correlation and $p$-value over training steps between the predicted prompt difficulty from our Bayesian surrogate and the empirical success rate. The strong correlation indicates that our method effectively predicts prompt difficulty without incurring costly LLM inferences.
1 Introduction
@@
Online Prompt Selection Matters in RL Finetuning: In RL finetuning of LLMs, a common operation is to execute random sampling from the prompt dataset, which serves for the chain-of-thought generation and policy optimization. However, random selection tends to suffer from efficiency and performance issues since it hardly circumvents sampling redundancies and fails to capture informative prompts critical for effective optimization while token generation itself occupies massive computation and memory resources (Zheng et al., 2025). Motivated by the importance of data quality in LLM training (Guo et al., 2025; Grattafiori et al., 2024), recent efforts have explored online prompt selection (Cui et al., 2025; Yang et al., 2024b; Yu et al., 2025; Bae et al., 2025; Meng et al., 2025; Zhang et al., 2025; Xiong et al., 2025). Technically, they dynamically curate training batches by prioritizing some prompt subset under certain criteria, e.g., filtering out low-quality prompts (Cui et al., 2025; Yu et al., 2025; Meng et al., 2025) or emphasizing moderately difficult ones, with difficulty quantified via success rate (Bae et al., 2025). These approaches exhibit superior performance and even accelerate the training process in certain scenarios. However, massive LLM calls for evaluating the difficulty of a large pool of candidate prompts inevitably incur substantial computational and memory overhead, raising severe concerns about efficiency (Chen et al., 2025).
Figure 2: Performance and computational efficiency of different prompt selection methods on Countdown. Our developed MoPPS surpasses uniform selection in both training efficiency and performance, while reducing computation by 79% fewer rollouts compared to DS (Yu et al., 2025).
@@
  2. 2.
    How can predicted outcome serve data selection for enhancing LLMs’ reasoning power?
Approximate Inference towards Prompt Difficulty for Active Selection: In response to these RQs, this work develops the Model Predictive Prompt Selection (MoPPS) method for online scoring the prompt difficulty in an amortized manner, which is simple to implement yet significantly improves learning efficiency in RL finetuning. Here, we formulate online prompt selection as a sequential decision-making problem and solve it with a dynamic Bernoulli bandit machine (Berry, 1972; Russo & Van Roy, 2014). In other words, each prompt is treated as an arm with stochastic binary rewards drawn from a latent variable as the success rate, and then we adopt posterior sampling to screen prompts in a streaming way. The lightweight sampling outcome of latent variables avoids exact evaluation of prompts, facilitates exploration as stochastic optimism, and supports informative prompt selection without extra LLM inference.
@@
  3. 3.
    Our framework is easy to implement and can be seamlessly integrated into a range of RL algorithms with various LLM backbones, complementing the pipeline of active RL finetuning.
Extensive experiments on complicated reasoning tasks spanning mathematics, planning, and vision-based geometry positively answer two RQs and reveal that MoPPS reliably predicts prompt difficulty, exhibiting high correlations with ground-truth prompt evaluation. Benefiting such predictability, our approach significantly accelerates RL finetuning, e.g., achieving $\bm{1.8\times}$ speedup over uniform sampling on Countdown (Pan et al., 2025), and yields better final performance, with over $\bm{24.4\%}$ relative improvements on the AIME24 dataset while training on MATH (Hendrycks et al., 2021).
@@ 2.1 Notations
$r_{\tau}^{t,j}\sim\mathrm{Bernoulli}(\gamma^{t}_{\tau}),\quad r_{\tau}^{t,j}=\begin{cases}1,&\text{if response $j$ is correct},\\ 0,&\text{otherwise},\end{cases}\quad j=1,\dots,k.$
@@
$p(r_{\tau}^{t,i})=(\gamma^{t}_{\tau})^{[r_{\tau}^{t,i}=1]}\cdot(1-\gamma^{t}_{\tau})^{[r_{\tau}^{t,i}=0]}\Rightarrow p(\bm{r}^{t}_{\tau}\mid\gamma^{t}_{\tau})=\binom{k}{s^{t}_{\tau}}\cdot(\gamma^{t}_{\tau})^{s^{t}_{\tau}}\cdot(1-\gamma^{t}_{\tau})^{k-s^{t}_{\tau}}\ \text{with}\ s^{t}_{\tau}\triangleq\sum_{j=1}^{k}r_{\tau}^{t,j}.$ (1)
For simplicity, this work focuses on binary reward signals in RL finetuning. However, the proposed method readily applies to richer reward forms such as format rewards (Pan et al., 2025), either by modeling them directly or by binarizing the signal through thresholding or rounding.
@@ 2.2 RL Finetuning for LLM
$\max_{\bm{\theta}}\;\mathbb{E}_{\tau\sim\mathcal{T},\;y\sim\pi_{\bm{\theta}}(\cdot|\tau)}\left[r(\tau,y)\right],$ (2)
@@
Proximal Policy Optimization (PPO).
@@
$\mathcal{J}_{\text{PPO}}(\bm{\theta})=\mathbb{E}_{\tau\sim\mathcal{T}_{t}^{\mathcal{B}},\;y_{\leq t}\sim\pi_{\bm{\theta}_{\text{old}}}(\cdot|\tau)}\left[\min\left(\rho_{t}(\bm{\theta})\cdot\hat{A}_{t},\;\text{clip}(\rho_{t}(\bm{\theta}),1-\epsilon,1+\epsilon)\cdot\hat{A}_{t}\right)\right],$ (3)
@@
Group Relative Policy Optimization (GRPO).
@@
$\mathcal{J}_{\text{GRPO}}$ $(\bm{\theta})=\mathbb{E}_{\tau\sim\mathcal{T}_{t}^{\mathcal{B}},\;\{y_{\tau}^{i}\}_{i=1}^{k}\sim\pi_{\bm{\theta}_{\text{old}}}(\cdot|\tau)}$ (4)
$\left[\frac{1}{k}\sum_{i=1}^{k}\frac{1}{|y_{\tau}^{i}|}\sum_{t=1}^{|y_{\tau}^{i}|}\left(\min\left(\rho_{i,t}(\bm{\theta})\cdot\hat{A}_{i},\;\text{clip}(\rho_{i,t}(\bm{\theta}),1-\epsilon,1+\epsilon)\cdot\hat{A}_{i}\right)-\beta D_{KL}(\pi_{\bm{\theta}}||\pi_{\text{ref}}\right)\right],$
@@
$\hat{A}_{i}=\frac{r_{\tau}^{i}-\text{mean}(\{r_{\tau}^{i}\}_{i=1}^{k})}{\text{std}(\{r_{\tau}^{i}\}_{i=1}^{k}}.$ (5)
Online Prompt Selection.
@@
$\mathcal{T}_{t}^{\mathcal{B}}=\left\{\tau\in\mathcal{T}_{t}^{\hat{\mathcal{B}}}\;\middle|\;0<\sum_{i=1}^{k}r^{i}_{\tau}<k\;\text{or}\;\text{std}\left(\{r^{i}_{\tau}\}_{i=1}^{k}\right)>0\right\}.$ (6)
@@
2.3 Model Predictive Task Sampling
MPTS (Wang et al., 2025a) is a recently proposed technique that amortizes expensive policy or data evaluation via active inference. It views the optimization process in task episodic learning as a generative process. With the help of streaming variational inference (Broderick et al., 2013; Nguyen et al., 2017), the risk predictive model $p(\ell|\tau,H_{t};\bm{\theta}_{t})$ is constructed to forecast policy evaluation outcomes or LLM training losses, where $\tau$ denotes a task identifier and $\ell$ is the associated evaluation metric, such as the negative average return obtained by policy $\bm{\theta}_{t}$ in MDP $\tau$ (Qu et al., 2025). The predictions are then combined with flexible acquisition criteria such as Upper Confidence Bound (UCB) (Auer, 2002) or posterior sampling (Qu et al., 2025) to achieve active sampling purposes. This way avoids additional expensive agent-environment interactions in RL or inference calls in LLMs.
In robust fast adaptation (Wang et al., 2023; Lv et al., 2024; Wang et al., 2024), task evaluation is crucial in subset selection, but incurs computational and annotation costs. MPTS has demonstrated its computational and sample efficiency in robust adaptive decision-making (Qu et al., 2025) and supervised finetuning (Wang et al., 2025a). However, it is designed for problems with continuous task spaces and task identifiers. In contrast, RL finetuning of LLMs involves discrete, language-defined tasks, without explicit task representations. Moreover, instead of robust adaptation, this domain prioritizes training efficiency due to the high computation, memory, and annotation costs, with the prompt success rate as an underexplored prediction target. Thus, distinguished from MPTS, this work focuses on (i) enabling predict-then-optimize in RL finetuning, i.e., amortizing the evaluation cost in the prompt selection, and (ii) identifying suitable acquisition criteria to accelerate training.
3 Method
This section surrounds technical solutions to two RQs in Sec. 1 and presents a principled and efficient framework for accelerating RL finetuning by model predictive prompt selection. First, we recast RL finetuning as a generative process and provide a corresponding probabilistic graphical model, where the success rate serves as the latent variable. To amortize the process of online prompt evaluation, we construct a collection of Bernoulli bandits and apply Bayesian inference to estimate the prompt-specific success rate. Then, we adopt Thompson sampling to obtain the predicted outcome and combine it with specific data selection criteria for efficient RL finetuning.
3.1 RL Finetuning as A Generative Process
@@
$p\bigl(\bm{\theta}_{0:T},\mathcal{T}_{0:T-1}^{\mathcal{B}},\mathcal{R}_{0:T-1}^{\mathcal{B}}\bigr)$ (7)
$\quad=p(\bm{\theta}_{0})\prod_{t=0}^{T-1}\;\underbrace{p\bigl(\mathcal{T}_{t}^{\mathcal{B}}\mid\bm{\theta}_{t}\bigr)}_{\text{{Prompt Selection}}}\;\underbrace{p\bigl(\bm{\theta}_{t+1}\mid\bm{\theta}_{t},\mathcal{R}_{t}^{\mathcal{B}},\mathcal{T}_{t}^{\mathcal{B}}\bigr)}_{\text{{Policy Optimization}}}\underbrace{\int p\bigl(\Gamma_{t}^{\mathcal{B}}\mid\mathcal{T}_{t}^{\mathcal{B}},\bm{\theta}_{t}\bigr)\;p\bigl(\mathcal{R}_{t}^{\mathcal{B}}\mid\Gamma_{t}^{\mathcal{B}}\bigr)\;d\Gamma_{t}^{\mathcal{B}}}_{\textbf{Prompt Evaluation}},$
@@
Figure 3: Probabilistic graphical model for RL finetuning of LLMs. The reward signal $\bm{r}^{t}_{\tau_{t,i}}$ is a set of binary values evaluating the $k$ generated responses, governed by the latent success rate $\gamma^{t}_{\tau_{t,i}}$. The prompt batch $\{\tau_{t,i}\}_{i=1}^{\mathcal{B}}$ is selected under specific criteria based on current LLM $\bm{\theta}_{t}$. The white and grey nodes respectively denote observed and latent variables.
@@
$\text{Online Prompt Selection:}\quad\mathcal{T}_{t}^{\hat{\mathcal{B}}}\xrightarrow{\text{{Evaluate}}}\{\mathcal{T}_{t}^{\hat{\mathcal{B}}},\mathcal{R}_{t}^{\hat{\mathcal{B}}}\}\xrightarrow{\text{{Filter}}}\{\mathcal{T}_{t}^{\mathcal{B}},\mathcal{R}_{t}^{\mathcal{B}}\}.$ (8)
@@
[…]
@@
3.2 Bayesian Inference towards Prompt Success Rate

### v5

@@
Abstract.
Recent advances have witnessed the effectiveness of reinforcement learning (RL) finetuning in enhancing the reasoning capabilities of large language models (LLMs). The optimization process often requires numerous iterations to achieve satisfactory performance, resulting in high computational costs due to the need for frequent prompt evaluations under intensive LLM interactions and repeated policy updates. Appropriate online prompt selection methods reduce iteration steps by prioritizing informative prompts during training, while the pipeline’s reliance on exhaustive prompt evaluation and subset selection for optimization still incurs substantial computational overhead due to frequent LLM inference calls. Distinguished from these direct evaluate-then-select schemes, this work investigates iterative approximate evaluation for arbitrary prompts and introduces Model Predictive Prompt Selection (MoPPS), a Bayesian risk-predictive framework that online estimates prompt difficulty without requiring costly LLM interactions. Technically, MoPPS models each prompt’s success rate as a latent variable, performs streaming Bayesian inference, and employs posterior sampling in a constructed multi-armed bandit machine, enabling efficient and adaptive prompt selection. Extensive experiments across mathematics, planning, and vision-based geometry tasks show that MoPPS reliably predicts prompt difficulty and accelerates training with significantly reduced LLM rollouts. Our code is available at https://github.com/thu-rllab/MoPPS.
Large Language Model, Reinforcement Learning, Reasoning Model, Online Prompt Selection, Active Data Sampling
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 9–13, 2025; Jeju Island, Republic of Korea.booktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD 2026), August 9–13, 2025, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780263isbn: 979-8-4007-2258-5/2026/08ccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Online Prompt Selection Matters in RL Finetuning: In RL finetuning of LLMs, random sampling from the prompt dataset is common for chain-of-thought generation and policy optimization. However, it often fails to capture informative prompts and suffers from inefficiency and redundancy, while token generation itself is resource-intensive (Zheng et al., 2025). Recent work emphasizes data quality (Guo et al., 2025; Grattafiori et al., 2024) and explores online prompt selection (Cui et al., 2025; Yang et al., 2024b; Yu et al., 2025; Bae et al., 2025; Meng et al., 2025; Zhang et al., 2025; Xiong et al., 2025), where training batches are curated by prioritizing prompts based on quality or difficulty (Cui et al., 2025; Yu et al., 2025; Meng et al., 2025; Bae et al., 2025). While these methods improve performance and even accelerate training, evaluating prompt difficulty across large candidate pools still incurs heavy computational overhead (Chen et al., 2025).
Figure 1. Performance and efficiency of prompt selection on Countdown. MoPPS outperforms uniform selection in training efficiency and performance and reduces rollouts by 78% compared to DS (Yu et al., 2025).
@@
  2. (2)
    How can predicted outcome serve active data sampling for enhancing LLMs’ reasoning power?
Approximate Inference towards Prompt Difficulty for Active Selection: In response to these RQs, this work develops the Model Predictive Prompt Selection (MoPPS) method for online scoring the prompt difficulty approximately, which is simple to implement yet significantly improves learning efficiency in RL finetuning. Here, we formulate online prompt selection as a sequential decision-making problem and solve it with a dynamic Bernoulli bandit-based data mining strategy (Berry, 1972; Russo and Van Roy, 2014). In other words, each prompt is treated as an arm with stochastic binary rewards drawn from a latent variable as the success rate, and then we adopt posterior sampling to screen prompts in a streaming way. The sampling outcome of latent variables avoids exact evaluation of prompts, facilitates exploration as stochastic optimism, and supports informative prompt selection without extra LLM inference.
@@
  3. (3)
    Our framework is easy to implement and can be seamlessly integrated into a range of RL algorithms with various LLM backbones, providing a system-efficient component that complements the pipeline of active RL finetuning.
Extensive experiments on complicated reasoning tasks spanning mathematics, planning, and vision-based geometry positively answer two RQs and reveal that MoPPS reliably predicts prompt difficulty, exhibiting high correlations with ground-truth evaluation. Benefiting such predictability, MoPPS significantly accelerates RL finetuning, e.g., achieving $\bm{1.8\times}$ speedup over uniform sampling on Countdown (Pan et al., 2025), and yields better performance, with over $\bm{24.4\%}$ relative improvements on the AIME24 while training on MATH (Hendrycks et al., 2021).
@@ 2.1. Notations
$\small r_{\tau}^{t,j}\sim\mathrm{Bernoulli}(\gamma^{t}_{\tau}),\ r_{\tau}^{t,j}=\begin{cases}1,&\text{if response $j$ is correct},\\ 0,&\text{otherwise},\end{cases}\ j=1,\dots,k.$
@@
(1) $p(r_{\tau}^{t,i})=(\gamma^{t}_{\tau})^{[r_{\tau}^{t,i}=1]}\cdot(1-\gamma^{t}_{\tau})^{[r_{\tau}^{t,i}=0]}$
$\Rightarrow p(\bm{r}^{t}_{\tau}\mid\gamma^{t}_{\tau})=\binom{k}{s^{t}_{\tau}}\cdot(\gamma^{t}_{\tau})^{s^{t}_{\tau}}\cdot(1-\gamma^{t}_{\tau})^{k-s^{t}_{\tau}}\ \text{with}\ s^{t}_{\tau}\triangleq\sum_{j=1}^{k}r_{\tau}^{t,j}.$
For simplicity, this work focuses on binary reward signals in RL finetuning. However, the proposed method readily applies to richer reward forms such as format rewards (Pan et al., 2025), either by modeling them directly or by binarizing them through thresholding or rounding.
@@ 2.2. RL Finetuning for LLM
(2) $\max_{\bm{\theta}}\;\mathbb{E}_{\tau\sim\mathcal{T},\;y\sim\pi_{\bm{\theta}}(\cdot|\tau)}\left[r(\tau,y)\right],$
@@
Proximal Policy Optimization (PPO)
@@
(3) $\mathcal{J}_{\text{PPO}}(\bm{\theta})$ $=\mathbb{E}_{\tau\sim\mathcal{T}_{t}^{\mathcal{B}},\;y_{\leq t}\sim\pi_{\bm{\theta}_{\text{old}}}(\cdot|\tau)}$
$\left[\min\left(\rho_{t}(\bm{\theta})\cdot\hat{A}_{t},\;\text{clip}(\rho_{t}(\bm{\theta}),1-\epsilon,1+\epsilon)\cdot\hat{A}_{t}\right)\right],$
@@
Group Relative Policy Optimization (GRPO)
@@
[…]
$\left.\left.\left.\rho_{i,t}(\bm{\theta})\cdot\hat{A}_{i},\;\text{clip}(\rho_{i,t}(\bm{\theta}),1-\epsilon,1+\epsilon)\cdot\hat{A}_{i}\right)-\beta D_{KL}(\pi_{\bm{\theta}}||\pi_{\text{ref}})\right)\right]$
@@
(5) $\hat{A}_{i}=\frac{r_{\tau}^{i}-\text{mean}(\{r_{\tau}^{i}\}_{i=1}^{k})}{\text{std}(\{r_{\tau}^{i}\}_{i=1}^{k}}.$
Online Prompt Selection
@@
(6) $\mathcal{T}_{t}^{\mathcal{B}}=\left\{\tau\in\mathcal{T}_{t}^{\hat{\mathcal{B}}}\;\middle|\;0<\sum_{i=1}^{k}r^{i}_{\tau}<k\;\text{or}\;\text{std}\left(\{r^{i}_{\tau}\}_{i=1}^{k}\right)>0\right\}.$
@@
2.3. Model Predictive Task Sampling
MPTS (Wang et al., 2025a) amortizes costly policy or data evaluation through active inference by modeling task optimization as a generative process. Using streaming variational inference (Broderick et al., 2013; Nguyen et al., 2017), it builds a predictive model $p(\ell|\tau,H_{t};\bm{\theta}_{t})$ to estimate evaluation metrics like returns in RL or training loss. These predictions guide active sampling via criteria such as UCB (Auer, 2002) or posterior sampling (Qu et al., 2025), reducing costly environment interactions or inference calls.
MPTS has shown its computational and annotation efficiency in adaptive decision-making (Qu et al., 2025) and supervised finetuning (Wang et al., 2025a), but assumes continuous task spaces with explicit identifiers. In particular, RL finetuning of LLMs involves discrete, prompt-defined tasks and emphasizes training efficiency. This work instead targets (i) amortizing prompt evaluation costs and (ii) identifying acquisition strategies to accelerate RL finetuning.
3. Method
This section surrounds two RQs from Sec. 1 and presents a principled framework for accelerating RL finetuning via model predictive prompt selection. We recast RL finetuning as a generative process with success rate as a latent variable, estimate prompt-specific performance using Bayesian Bernoulli bandits, and adopt Thompson sampling with selection criteria for efficient RL finetuning.
Figure 2. Probabilistic graphical model for RL finetuning of LLMs. The reward signal $\bm{r}^{t}_{\tau_{t,i}}$ is a set of binary values evaluating the $k$ generated responses, governed by the latent success rate $\gamma^{t}_{\tau_{t,i}}$. The prompt batch $\{\tau_{t,i}\}_{i=1}^{\mathcal{B}}$ is selected under specific criteria based on current LLM $\bm{\theta}_{t}$. The white and grey nodes respectively denote observed and latent variables.
3.1. RL Finetuning as A Generative Process
@@
(7) $p\bigl(\bm{\theta}_{0:T},\mathcal{T}_{0:T-1}^{\mathcal{B}},\mathcal{R}_{0:T-1}^{\mathcal{B}}\bigr)\quad=\quad p(\bm{\theta}_{0})\prod_{t=0}^{T-1}\;\underbrace{p\bigl(\mathcal{T}_{t}^{\mathcal{B}}\mid\bm{\theta}_{t}\bigr)}_{\text{{Prompt Selection}}}\;$
$\underbrace{p\bigl(\bm{\theta}_{t+1}\mid\bm{\theta}_{t},\mathcal{R}_{t}^{\mathcal{B}},\mathcal{T}_{t}^{\mathcal{B}}\bigr)}_{\text{{Policy Optimization}}}\underbrace{\int p\bigl(\Gamma_{t}^{\mathcal{B}}\mid\mathcal{T}_{t}^{\mathcal{B}},\bm{\theta}_{t}\bigr)\;p\bigl(\mathcal{R}_{t}^{\mathcal{B}}\mid\Gamma_{t}^{\mathcal{B}}\bigr)\;d\Gamma_{t}^{\mathcal{B}}}_{\textbf{Prompt Evaluation}},$
@@
(8) Online Prompt Selection:
$\mathcal{T}_{t}^{\hat{\mathcal{B}}}\xrightarrow{\text{{Evaluate}}}\{\mathcal{T}_{t}^{\hat{\mathcal{B}}},\mathcal{R}_{t}^{\hat{\mathcal{B}}}\}\xrightarrow{\text{{Filter}}}\{\mathcal{T}_{t}^{\mathcal{B}},\mathcal{R}_{t}^{\mathcal{B}}\}.$
@@
(9) $p(\mathcal{T}_{t}^{\mathcal{B}}\mid\bm{\theta}_{t})$ $=\int p(\mathcal{T}_{t}^{\mathcal{B}}\mid\mathcal{R}_{t}^{\hat{\mathcal{B}}},\mathcal{T}_{t}^{\hat{\mathcal{B}}})p(\mathcal{T}_{t}^{\hat{\mathcal{B}}})$
$\underbrace{\int p(\Gamma_{t}^{\hat{\mathcal{B}}}\mid\mathcal{T}_{t}^{\hat{\mathcal{B}}},\bm{\theta}_{t})\,p(\mathcal{R}_{t}^{\hat{\mathcal{B}}}\mid\Gamma_{t}^{\hat{\mathcal{B}}})\,d\Gamma_{t}^{\hat{\mathcal{B}}}}_{\text{{Extra Prompt Evaluation}}}\;d\mathcal{R}_{t}^{\hat{\mathcal{B}}}\;d\mathcal{T}_{t}^{\hat{\mathcal{B}}},$
@@
3.2. Bayesian Inference towards Prompt Success Rate